A mutant dwarf tomato plant with yellow fruit is self-fertilized and the following progeny are produced:

Normal size, yellow fruit: 39 plants

Normal size, red fruit: 13 plants

Dwarf, yellow: 111 plants

Dwarf, red : 35 plants

a. What are the phenotypes, alleles, and possible genotypes for the two traits? For the alleles designations, you must use the common convention for assigning alleles for plants, includes symbols consistent with how alleles interact.

b. What is the most likely genotype of the plant that was self-fertilized?
Answer #1

Answer:

Based on the information, dwarf (D) is dominant over Norma size (d) and yellow (Y) is dominant over red (y).

As all traits are produced and numbers of alll phenotypes are more or less resemble to 9:3:3:1. Hence, the genotype of the plant that was self-fertilized = DdYy

DdYy x DdYy ----- parents

Progeny;

D_Y_= 9/16= dwarf, yellow

D_yy = 3/16= dwarf, red

ddY- = 3/16= normal size, yellow

ddyy = 1/16= normal size, red